Check the condition of integrability & solve

yzdx + 2xzdy − 3xydz = 0.
1
Expert's answer
2020-11-23T18:55:21-0500

Let us denote "U=yz\cdot i + 2xz\cdot j − 3xy\cdot k." Then


"curl\ U=\left|\begin{array}{ccc} i & j & k \\ \frac{\partial }{\partial x} & \frac{\partial }{\partial y} & \frac{\partial }{\partial z}\\ yz & 2xz & − 3xy \end{array}\right|=(-3x-2x)i-(-3y-y)j+(2z-z)k="

"=-5x\cdot i +4y\cdot j+z\cdot k"


Since the scalar product "U\cdot curl U=(yz\cdot i + 2xz\cdot j − 3xy\cdot k)\cdot(-5x\cdot i +4y\cdot j+z\cdot k)="


"-5xyz+8xyz-3xyz=0", we conclude that "yzdx + 2xzdy \u2212 3xydz = 0" is integrable.


Let us divide both part of the equation "yzdx + 2xzdy \u2212 3xydz = 0" by "xyz":


"\frac{dx}{x}+2\frac{dy}{y}-3\frac{dz}{z}=0"


It folows that "\int\frac{dx}{x}+2\int\frac{dy}{y}-3\int\frac{dz}{z}=0" and therefore, "\ln |x|+2\ln|y|-3\ln|z|=\ln|C|".

Then "\ln|xy^2|=\ln|Cz^3|". Consequently, the solution is the following:


"xy^2=Cz^3"